[Ncep.nhc.nco_contacts] TSB Morning Rounds - Tue Jul 11, 2017

TSB Admin nhc.tsbadmin at noaa.gov
Tue Jul 11 16:25:52 UTC 2017


Here is today's summary of NHC computer operations:
--- Craig Mattocks

----------

TAFB:
1. Multiple problems with AWIPS-II collaboration software. Update from TSB  
AWIPS-II Sysadmin Chris Mello: This appears to be an issue in the baseline  
software. Other sites are suffering from the same problems. The NCF ticket  
(142579) for this issue has been elevated.

2. Bad connection to the large AWIPS-II/GFE monitor at the TAFB Pacific  
Desk. Update: Chris Mello replaced the monitor cable, which seems to have  
fixed the problem.

3. Corrupted TAFB forecaster accounts in AWIPS-II. Update: TSB AWIPS-II  
Sysadmin Chris Mello has rebuilt the damaged accounts.
HSU:
1. ASCAT satellite scatterometer data stopped coming into N-AWIPS last  
evening. The Hurricane Specialists (and OPC) called the SDM (Senior Duty  
Meteorologist), who called the Earth System Prediction Capability (ESPC) at  
NESDIS. They were aware of the problem and had a programmer investigating  
the outage. NCO's Big Brother alerts cleared 77 minutes later and data  
began flowing again during the midnight shift.

2. The Hurricane Specialists are unable to print from the Atlantic N-AWIPS  
workstations Tilefish & Ridges. Chris Mello submitted a Priority Level 1 -  
Immediate (ASAP/1 hour) NCO Help Desk ticket.

3. Incorrect date for the 09Z issuance of the Time-of-Arrival (ToA)  
graphic. Update: TSB Developer Matt Onderlinde has corrected the PDT time  
zone offset problem.
Other items:
1. Late arriving GFS model products. After the 6-day outage due to the Cray  
GPFS filesystem split, the amount of data crossing the network between the  
WCOSS supercomputing clusters in Orlando (Surge/Gyre) and their  
counterparts in Reston (Luna/Tide) overwhelmed the operational dataflow.  
NCO immediately closed all development transfer queues and halted model  
data (com, nwges, etc.) catchup efforts, waiting for DBNet to stabilize.  
Once dataflow is able to catch up, NCO will resume backfilling model data  
but will continue to keep development transfers turned off. The current  
expectation is that the data won't be fully caught up until at least  
tomorrow (Wednesday) morning.
NHC RFCs:


-------------- next part --------------
An HTML attachment was scrubbed...
URL: https://www.lstsrv.ncep.noaa.gov/pipermail/ncep.nhc.nco_contacts/attachments/20170711/bebb9bd5/attachment.html 


More information about the Ncep.nhc.nco_contacts mailing list